25. Today (1-0), an investor purchased a 20 year bond with a 5.00% coupon and a face value of $100,000 for $106,550. In six months (T 0.5) interest rates have decreased by 0.50% and the investor decides to sel the bond immediately after receiving the first coupon payment. What is the investors total gain (loss) on the bond? HINT: Total Gain (Loss) = Price Change in Bond + Coupon A. ($6,548) B. ($6,048) C. $7,130 D. $7,602 E. $7,630 Assume all future cash flows (FCFF, FCFE, and Dividends) will grow at a constant sustainable growth rate (g) in perpetuity; note this growth rate is dependent on a companys return on equity and dividend payout policy. The companys cost of debt is 20% while the companys cost of equity is 30%; the company has an effective tax rate of 35%. Use the following information in the table below to help answer problems 26-27: FCFE Today (T 0) FCFF Today (7-0) Shareholder Equity Total Debt Total Assets Net Income Dividends Shares Outstanding in millions 800 750 400 600 1000 100 40 25 26. An activist investor who wants to purchase all the companys shares would be willing to pay approximately. Do not use the DDM. Use either FCFF or FCFE, whichever is appropriate a. $165 per share b. $176 per share c. $213 per share d. $230 per share e. $245 per share 27. A private equity group who wants to purchase all of the companys assets would be willing to pay approximately. Do not use the DDM. Use either FCFF or FCFE, whichever is appropriate a. $9,583M b. $10,222M c. $10,518M d. $17,969M e. $19,167M
Answer #1

Since you have asked multiple unrelated questions, I will address the first one.

Price of the bond, PV = 106,550

Face Value of the bond, FV = 100,000

Annual coupon = 5%

Frequency = Semi annual

Coupon per period = 5% / 2 = 2.5% x 100,000 = 2,500 = Payment

Time to maturity = 20 years

Periods to maturity = Nos. of half years in 20 years = 2 x 20 = 40

Let's assume that YTM of this bond is y. Yield per period = y / 2

Yield can be calculated using RATE function of excel.

Yield per period = y / 2 = RATE(Period, Payment, PV, FV) = RATE(40, 2500, -106550, 100000) = 2.25%

Hence, yield = y = 2 x 2.25% = 4.50%

6 months down the line,

Yield = y = 0.5% reduction from previous yield = 4.50% - 0.5% = 4.00%

Yield per period = y/2 = 4.00% / 2 = 2.00%

Period to maturity, N = 19.5 years = 2 x 19.5 = 39 half years

Coupon payment per period, C = 2,500

Face Value, FV = 100,000

Hence price of bond = PV of future coupon periods + PV of FV

=\frac{C}{\frac{y}{2}}\times (1- (1+y/2)^{-N}) + FV\times(1+y/2)^{-N}

= 2500 / 0.02 x [1 - (1.02)(-39)] + 100,000 x (1.02)(-39)

= $ 113,453

Total gain = price change in bond + coupon received = 113,453 - 106,550 + 2,500 = 9,403

None of the options contain this as an answer. Hence, none of the options are correct.

Please check the question and the options at your end.